# Noise Synthesis for Low-Light Image Denoising with Diffusion Models

### v1

@@ Abstract
Low-light photography produces images with low signal-to-noise ratios due to limited photons. In such conditions, common approximations like the Gaussian noise model fall short, and many denoising techniques fail to remove noise effectively. Although deep-learning methods perform well, they require large datasets of paired images that are impractical to acquire. As a remedy, synthesizing realistic low-light noise has gained significant attention. In this paper, we investigate the ability of diffusion models to capture the complex distribution of low-light noise. We show that a naive application of conventional diffusion models is inadequate for this task and propose three key adaptations that enable high-precision noise generation without calibration or post-processing: a two-branch architecture to better model signal-dependent and signal-independent noise, the incorporation of positional information to capture fixed-pattern noise, and a tailored diffusion noise schedule. Consequently, our model enables the generation of large datasets for training low-light denoising networks, leading to state-of-the-art performance. Through comprehensive analysis, including statistical evaluation and noise decomposition, we provide deeper insights into the characteristics of the generated data.
@@ 1 Introduction
Figure 1: Denoising results from networks trained with synthetic data generated by different synthesis methods, with a network trained on paired real data included for reference. The sample image is from the ELD test set [52] (cropped for visualization). The effectiveness of trained denoiser strongly depends on the quality of the noise data used for training. Our approach is to utilize diffusion models to generate synthetic noise data that closely fits real situations in low-light scenarios, enabling the trained model to denoise images effectively. (Best viewed when zoomed in.)
Obtaining noiseless photographs in low-light settings is extremely challenging. Although extending the exposure time can gather more light and enhance brightness in still scenes, this solution falls short for dynamic scenes due to the necessity of a static environment during image capture. Alternatively, increasing the system gain expands the RAW signal’s dynamic range, but this operation also dramatically increases the noise level in the captured image.
@@
In contrast, diffusion models [25, 14, 42] have shown strong performance in capturing complex multi-modal data distributions. In this paper, we investigate their application to the task of low-light noise synthesis. While a direct application of conventional diffusion models proves inadequate for this task, our proposed framework achieves state-of-the-art noise synthesis performance thanks to several adaptations. We discuss the role of each adaptation by carefully analyzing various statistics of the generated noise in Sec. 4.4. Our contributions are as follows: (1) We demonstrate that a two-branch neural network is an effective approach for modeling signal-dependent and signal-independent noise. (2) We incorporate positional encoding to account for the spatially varying nature of signal-independent noise. (3) We show empirically and theoretically that the diffusion noise schedule is critical for accurately capturing the noise variance. (4) We obtain state-of-the-art low-light noise synthesis performance, leading to better denoising performance compared to models trained on data synthesized by other techniques, as illustrated in Fig. 1.
@@ 2 Related Work
This model remains unable to capture all relevant noise sources. Subsequent works [48, 12, 52, 57, 20] introduce more complex and accurate noise models. In addition to the previously mentioned noise sources, these models incorporate other components of real-world noise, such as fixed-pattern noise and color-biased read noise. Recently, the ELD paper [52] introduces a noise model that accounts for multiple noise sources. After calibration, this noise model was utilized to synthesize noisy data for training an image-denoising CNN. Despite these advancements in physical noise modeling, a discrepancy persists between synthetic and real-world noise. In particular, modeling the signal-independent component of real noise is extremely challenging.
@@
Deep Learning-Based Noise Modeling. Another research line focuses on using deep neural networks to capture real-world noise more precisely. C2N [27] proposes to train a generative adversarial network (GAN) [23] with unpaired noisy and clean images for noisy image generation. The Noise Flow method [2] introduces the use of normalizing flows [15, 28, 29] to model real-world RAW noise. [30] also adopts normalizing flows for noise modeling, but in the output sRGB space.
@@
Normalizing flows, however, struggle to capture highly multimodal or intricate structures in the data, and performance can be sensitive to the choice of flow architecture, hence necessitating careful tuning and experimentation [6]. As for GANs, they are notorious for their unstable training process [24, 36]. Diffusion models have demonstrated strong capabilities in modeling complex data distributions [41, 42, 39]. Therefore, in this paper, we investigate the use of diffusion models to model low-light noise to enable the generation of more realistic data samples.
@@ 3 Method
Conventional diffusion models for image generation cannot be naively adapted to low-light noise generation. We propose several key adaptations that enable high-precision noise generation. (Please refer to the appendix for preliminaries on diffusion models.)
Figure 2: (a) The proposed diffusion framework for synthetic data generation. The forward process progressively destroys a real low-light noise $\mathbf{n}_{0}$ by adding Gaussian noise. In the reverse process, $\tilde{\mathbf{n}}_{T}$ is sampled from a Gaussian distribution and progressively denoised by a network back to the low-light noise $\tilde{\mathbf{n}}_{0}$. The network is conditioned on the clean image $\mathbf{x}$, pixel coordinates $\mathbf{c}$, and camera settings $\mathbf{z}$, as all these factors influence the noise pattern. Once the low-light noise $\tilde{\mathbf{n}}_{0}$ is obtained, it can be added to the clean image $\mathbf{x}$ to create the corresponding noisy image $\tilde{\mathbf{y}}$, thus forming a synthetic clean-noisy pair $(\mathbf{x},\tilde{\mathbf{y}})$. (b) The detailed architecture of the two-branch diffusion network, shown from left to right: the MLP branch, the UNet branch, and the architecture of the down-sampling block (with up-sampling blocks having a similar structure).
@@ 3.1 Framework for Low-light Noise Generation
The overall framework is depicted in Fig. 2 (a). The initial input for the forward process of the diffusion model is the low-light noise $\mathbf{n}_{0}$, obtained by subtracting the clean image from its noisy counterpart (which is obtained by multiplying the low-exposure image by the exposure ratio to match the dynamic range of the clean image). Random samples of Gaussian noise are then added to $\mathbf{n}_{0}$, gradually approximating a standard normal distribution after $T$ steps. For the backward process, a network is employed to predict the added Gaussian noise and denoise $\tilde{\mathbf{n}}_{T}$ back to $\tilde{\mathbf{n}}_{0}$.
As signal-dependent noise is correlated to the clean image and camera settings, we condition our diffusion network on these factors. We also encode pixel positions into the diffusion network to better capture spatially related noise components, such as fixed-pattern noise, which is salient in low-light images. Hence, the reverse process of our diffusion model is formulated as $\tilde{\mathbf{n}}_{t}=f_{\theta}(\tilde{\mathbf{n}}_{t-1},t,\mathbf{x},\mathbf{c},\mathbf{z})$, where $t$ is the diffusion timestep, $\mathbf{x}$ is the clean image, $\mathbf{c}$ are the pixel coordinates, and $\mathbf{z}$ is the camera setting (including ISO value and the exposure ratio). Conditioning the diffusion model on the clean image is done by concatenating it with the input noise image. Regarding the pixel coordinates, the camera settings, and the diffusion timestep, further explanation is provided in Sec. 3.2.
3.2 Network Architecture
Two-Branch Architecture. We design our diffusion network with two branches as shown in Fig. 2 (b): a Multilayer perceptron (MLP) branch and a UNet branch. This design addresses the distinct nature of different noise components. Noise components such as shot noise only depend on the current clean pixel value. Hence, their generation should not be influenced by neighboring pixels, making the MLP branch well-suited for handling these characteristics. Conversely, noise components that are locally correlated are better captured by the UNet architecture that excels at modeling local dependencies. In the experimental section, we will further validate the effectiveness of this design.
@@ 3.2 Network Architecture
Figure 3: Impact of noise schedules. (a) Visualization of the mean-variance correspondences of the real and generated Poisson noises across different reverse steps ($t=4,2,0$) under the cosine noise-scheduling scheme. (b) Different noise scheduling schemes, the y-axis denotes the coefficient $\sqrt{1-\overline{\alpha}_{t}}$ of the noise part in Eq. 3 of the appendix. (c) Results of different noise scheduling schemes, where sigmoid2 and sigmoid3 perform the best in our 1D toy example. Zoom in for a clearer view.
@@ 3.3 Variance Preserving Noise Schedule
As $\sigma_{1}^{2}$ increases, the variance of the synthesized noise decreases. Therefore, to preserve the variance of $U$, $\sigma_{1}^{2}$ should be minimized. In other words, in the final steps of the reverse process (where the diffusion model focuses more on high-frequency details [49]), the diffusion model should be exposed to minimal Gaussian noise; otherwise, the generated output will be overly smoothed. Next, we validate this hypothesis through experiments on a 1D toy example by testing different noise schedules.
Performance of Different Noise Schedules. For simplicity, we use a 1D Poisson distribution as a basic approximation of real-world noise. We train a 1D diffusion model using a 1D UNet to approach this distribution. The diffusion model employs the commonly used cosine noise schedule, which is plotted in Fig. 3 (b) in pink.
@@
Dataset Ratio Real-noise-based Physics-based DNN-based
Paired data Poisson-Gaussian ELD [52] NoiseFlow [2] LRD [56] Ours
Results without any post-processing
SID $\times$100 42.40 / 0.958 39.79 / 0.929 41.21 / 0.952 38.42 / 0.917 39.93 / 0.924 42.55 / 0.957
$\times$250 39.68 / 0.942 34.71 / 0.863 38.54 / 0.929 34.22 / 0.850 35.42 / 0.857 39.67 / 0.943
$\times$300 36.95/ 0.927 31.74 / 0.783 35.35 / 0.908 32.07 / 0.784 32.50 / 0.783 37.24 / 0.927
@@
$\times$200 41.04 / 0.895 39.00 / 0.831 43.20 / 0.954 39.35 / 0.871 40.99 / 0.879 41.36 / 0.912
Results post-processed by the illumination correction method used in the ELD paper
SID $\times$100 42.95 / 0.958 41.05 / 0.936 41.95 / 0.953 40.21 / 0.925 40.97 / 0.930 43.30 / 0.958
$\times$250 40.27 / 0.943 36.63 / 0.885 39.44 / 0.931 36.15 / 0.870 37.00 / 0.875 40.53 / 0.944
$\times$300 37.32 / 0.928 33.34 / 0.811 36.36 / 0.911 33.27 / 0.803 33.74 / 0.803 37.68 / 0.928
[…]
$\times$200 41.70 / 0.912 41.00 / 0.880 43.43 / 0.954 40.26 / 0.885 41.71 / 0.899 42.25 / 0.924
Table 1: Quantitative evaluation of the SID and ELD datasets in terms of PSNR and SSIM across various exposure ratios. Results are provided for two cases: one without any post-processing, and another post-processed using the illumination correction method from the ELD paper. Best results are highlighted in bold, and second-best results are underlined. Our method outperforms all other synthesis methods across most settings.
@@ 4.1 Evaluation Protocols
Noise Statistics. We also analyze the pixel-wise marginal distributions between the generated and real noise images in the test set, similar to the approach presented in the Noise Flow paper [2]. This evaluation prioritizes the model’s ability to capture the fundamental characteristics of the noise distribution. Specifically, we compute the histograms of real-world and generated noise and report the discrete KL divergence between these histograms.
@@ 4.2 Implementation Details
We train our diffusion model using the SID Sony RAW image training set. For evaluation, in line with the approach used in the ELD paper [52], we evaluate our denoising network on a combination of the SID test and validation sets (referred to as the SID test set hereafter), using the same evaluation images as in the ELD paper. We also evaluate our model on the ELD Sony dataset. Further implementation details are provided in the appendix.

### v2

# Dark Noise Diffusion: Noise Synthesis for Low-Light Image Denoising

@@ Abstract
Low-light photography produces images with low signal-to-noise ratios due to limited photons. In such conditions, common approximations like the Gaussian noise model fall short, and many denoising techniques fail to remove noise effectively. Although deep-learning methods perform well, they require large datasets of paired images that are impractical to acquire. As a remedy, synthesizing realistic low-light noise has gained significant attention. In this paper, we investigate the ability of diffusion models to capture the complex distribution of low-light noise. We show that a naive application of conventional diffusion models is inadequate for this task and propose three key adaptations that enable high-precision noise generation: a two-branch architecture to better model signal-dependent and signal-independent noise, the incorporation of positional information to capture fixed-pattern noise, and a tailored diffusion noise schedule. Consequently, our model enables the generation of large datasets for training low-light denoising networks, leading to state-of-the-art performance. Through comprehensive analysis, including statistical evaluation and noise decomposition, we provide deeper insights into the characteristics of the generated data.
Index Terms:
Computational Photography - Low-light Imaging - Image Denoising - Diffusion Models
@@ 1 Introduction
Obtaining noiseless photographs in low-light settings is extremely challenging. Although extending the exposure time enhances brightness by gathering more light, this solution falls short for dynamic scenes due to the necessity of a static environment during image capture. Alternatively, increasing the system gain expands the RAW signal’s dynamic range, but this operation also dramatically increases the noise level in the captured image.
@@
Figure 1: Examples of noise images generated by our Dark Noise Diffusion (NoiseDiff) method for three images from the SID Sony dataset [4]. Noise values are represented following a color map displayed on the right of the generated noise images. The generated noise images closely resemble real noise images. The distribution of the generated noise also aligns well with the real noise distributions.
@@
In contrast, diffusion models [16, 17, 18] have shown strong performance in capturing complex multi-modal data distributions. In this paper, we investigate their application to the task of low-light noise synthesis. While a direct application of conventional diffusion models proves inadequate for this task, our proposed framework, Dark Noise Diffusion (NoiseDiff), achieves state-of-the-art noise synthesis performance thanks to several adaptations, regarding the architecture of the diffusion network, and the diffusion noise scheduling scheme. As illustrated in Fig. 1, our method generates noise with high precision, effectively capturing both local and global features. For each proposed adaptation, we investigate its relevance by carefully analyzing various statistics of the generated noise in Sec. 4.5. Our contributions are as follows: (1) We demonstrate that a two-branch neural network is an effective approach for modeling signal-dependent and signal-independent noise. (2) We incorporate positional encoding to account for the spatially varying nature of signal-independent noise. (3) We show empirically and theoretically that the diffusion noise schedule is critical for accurately capturing the noise variance. (4) We obtain state-of-the-art low-light noise synthesis performance, leading to better denoising performance compared to models trained on data synthesized by other techniques.
@@ 2 Related Work
During photon collection, let $u^{*}$ denote the expected number of incident photons, $\alpha$ the quantum efficiency of the sensor, and $g$ the system gain of the analog amplifier, commonly referred to as ISO. The ideal signal, assuming perfect photon detection, is then given by: $x^{*}=g\alpha u^{*}$. However, the actual number of collected photons, $u$, is modeled as a Poisson distribution with mean $u^{*}$: $u\sim\mathcal{P}(u^{*})$. In addition to photon shot noise, the sensor introduces dark current noise, $n_{d}$, which is usually modeled as a Gaussian distribution: $n_{d}\sim\mathcal{N}(0,\sigma_{d}^{2})$.
Subsequently, during the signal readout process, additional signal-independent noise sources, including read noise, are introduced. These are commonly modeled as a zero-mean Gaussian distribution $n_{r}\sim\mathcal{N}(0,\sigma_{r}^{2})$. The actual observed signal $x$ can be formulated as
[…]
where $\sigma^{2}=g^{2}\sigma_{d}^{2}+\sigma_{r}^{2}$. The noise component, defined as the deviation from the ideal signal, is then
$n=x-x^{*}=g\alpha\mathcal{P}(u^{*})-g\alpha u^{*}+\mathcal{N}(0,\sigma^{2})\;.$ (2)
However, this model remains unable to capture all relevant noise sources. Subsequent works [5, 6, 9, 7, 8] introduce more complex and accurate noise models. In addition to the previously mentioned noise sources, these models incorporate other components of real-world noise, such as fixed-pattern noise and color-biased read noise. Recently, the ELD paper [9] introduces a noise model that accounts for other noise sources. After calibration, this noise model was utilized to synthesize noisy data for training an image-denoising CNN. Despite these advancements in physical noise modeling, a discrepancy persists between synthetic and real-world noise. In particular, modeling the signal-independent component of real noise is extremely challenging.
@@
Deep Learning-Based Noise Modeling. Another research direction focuses on using deep neural networks to capture real-world noise more precisely. C2N [10] proposes to train a generative adversarial network (GAN) [46] with unpaired noisy and clean images for noisy image generation. The Noise Flow method [13] introduces the use of normalizing flows [47, 48, 49] to model real-world RAW noise. [50] also adopts normalizing flows for noise modeling, but in the output sRGB space.
@@
Despite these advances, these methods suffer from several limitations. Normalizing flows struggle to capture multimodal or intricate structures in the data, and performance is sensitive to the choice of flow architecture, hence requiring careful tuning and experimentation [15]. As for GANs, they are notorious for their unstable training process [51, 52]. Diffusion models have demonstrated strong capabilities in modeling complex data distributions [53, 18, 54]. Therefore, in this paper, we investigate the use of diffusion models to model low-light noise to enable the generation of more realistic data samples.
@@ 3 Method
Conventional diffusion models for image generation cannot be directly applied to low-light noise generation. We propose several crucial adaptations that enable high-precision noise generation, regarding (1) the architecture of the diffusion network, (2) the conditioning of our generation algorithm, and (3) the noise schedule of the diffusion process. (Please refer to the appendix for preliminaries on diffusion models.)
Figure 2: (a) The proposed Dark Noise Diffusion (NoiseDiff) framework for synthetic data generation. The forward process progressively destroys a real low-light noise $\mathbf{n}_{0}$ by adding Gaussian noise. In the reverse process, $\tilde{\mathbf{n}}_{T}$ is sampled from a Gaussian distribution and progressively denoised by a network back to the low-light noise $\tilde{\mathbf{n}}_{0}$. The network is conditioned on the clean image $\mathbf{x}$, pixel coordinates $\mathbf{c}$, and camera settings $\mathbf{z}$, as all these factors influence the noise pattern. Once the low-light noise $\tilde{\mathbf{n}}_{0}$ is obtained, it can be added to the clean image $\mathbf{x}$ to create the corresponding noisy image $\tilde{\mathbf{y}}$, thus forming a synthetic clean-noisy pair $(\mathbf{x},\tilde{\mathbf{y}})$. (b) The detailed architecture of the two-branch diffusion network, shown from left to right: the MLP branch, the UNet branch, and the architecture of the down-sampling block (with up-sampling blocks having a similar structure).
@@ 3.1 Framework for Low-light Noise Generation
The overall NoiseDiff framework is depicted in Fig. 2 (a). The initial input for the forward process of the diffusion model is a low-light noise image $\mathbf{n}_{0}$, obtained by subtracting a clean image from its noisy counterpart (the latter is obtained by multiplying the low-exposure image by the exposure ratio to match the dynamic range of the clean image). Random samples of Gaussian noise are then added to $\mathbf{n}_{0}$, gradually approaching a standard normal distribution after $T$ steps. For the backward process, a network is employed to predict the added Gaussian noise and denoise $\tilde{\mathbf{n}}_{T}$ back to $\tilde{\mathbf{n}}_{0}$.
As signal-dependent noise is correlated to the clean image and camera settings, we condition our diffusion network on these factors. We also encode pixel positions into the diffusion network to better capture spatially related noise components, such as fixed-pattern noise, which is salient in low-light images. Hence, the reverse process of our diffusion model is formulated as $\tilde{\mathbf{n}}_{t}=f_{\theta}(\tilde{\mathbf{n}}_{t-1},t,\mathbf{x},\mathbf{c},\mathbf{z})$, where $t$ is the diffusion timestep, $\mathbf{x}$ is the clean image, $\mathbf{c}$ are the pixel coordinates, and $\mathbf{z}$ is the camera setting (including ISO value and the exposure ratio). Conditioning the diffusion model on the clean image is done by concatenating it with the input noise image. Regarding the encoding of the pixel coordinates, the camera settings, and the diffusion timestep, further explanation is provided in Sec. 3.2.
3.2 Diffusion Network Architecture
Two-Branch Architecture. We design our diffusion network with two branches as shown in Fig. 2 (b): a Multilayer perceptron (MLP) branch and a UNet branch. This design addresses the distinct nature of different noise components. Noise components such as shot noise only depend on the current clean pixel value. Hence, their generation should not be influenced by neighboring pixels, making the MLP branch well-suited for handling these characteristics. Conversely, noise components that are locally correlated are better captured by the UNet architecture that excels at modeling local dependencies. It is important to note that we do not explicitly enforce the MLP branch to model only shot noise or the UNet branch to model other noise components, as the ground-truth separation of these noise types is unavailable. However, as will be demonstrated in the experiment section, this design improves noise modeling accuracy.
@@ 3.2 Diffusion Network Architecture
Figure 3: Impact of noise schedules. (a) Visualization of the mean-variance correspondences of the real and generated Poisson noises across different reverse steps ($t=5\rightarrow 0$) under the cosine noise-scheduling scheme. (b) Different noise scheduling schemes, the y-axis denotes the coefficient $\sqrt{1-\overline{\alpha}_{t}}$ of the noise part in Eq. 5 of the appendix. (c) Results of different noise scheduling schemes, where sigmoid2 and sigmoid3 perform the best in our 1D toy example. Zoom in for a clearer view.
@@ 3.3 Variance Preserving Noise Schedule
As $\sigma_{1}^{2}$ increases, the variance of the synthesized noise decreases. Therefore, to preserve the variance of the signal $U$, $\sigma_{1}^{2}$ should be minimized. In other words, in the final steps of the reverse process (where the diffusion model focuses more on high-frequency details [59]), the diffusion model should be exposed to minimal Gaussian noise; otherwise, the generated output will be overly smoothed.
This theory can be easily extended to more complex noise distributions, as the same property holds for Gaussian Mixture Models (GMMs), which can approximate any distribution, given a sufficient number of components. The details can be found in the appendix.
Next, we validate this hypothesis through experiments on a 1D toy example by testing different noise schedules.
Performance of Different Noise Schedules. For simplicity, we use a 1D Poisson distribution as a basic approximation of real-world noise. We train a 1D diffusion model using a 1D UNet to approach this distribution. The diffusion model employs the commonly used cosine noise schedule, which is plotted in Fig. 3 (b) in magenta.
@@
In the appendix, we also conduct an experiment on a set of 1D Tukey-Lambda (TL) distributions to further validate this theory.
TABLE I: Denoising performance on the SID and ELD datasets in terms of PSNR and SSIM across various exposure ratios. Noise synthesis methods are divided into two groups: W/ Calibration, which use dark shading data while training the denoising network; and W/o Calibration, which do not use dark shading. Best results are highlighted in bold, and second-best results are underlined. Our NoiseDiff and NoiseDiff* outperform other noise synthesis methods in most cases under the same evaluation setting.
Dataset Ratio Real-noise-based W/o Calibration W/ Calibration
Real data Poisson-Gaussian ELD [9] Starlight [11] NoiseFlow [13] NoiseDiff PMN [60] LRD [12] NoiseDiff*
SID $\times$100 42.95 / 0.958 41.05 / 0.936 41.95 / 0.953 40.47 / 0.926 40.20 / 0.925 43.30 / 0.958 43.47 / 0.961 43.16 / 0.958 43.92 / 0.961
$\times$250 40.27 / 0.943 36.63 / 0.885 39.44 / 0.931 36.25 / 0.858 36.15 / 0.870 40.53 / 0.944 41.04 / 0.947 40.69 / 0.941 41.28 / 0.946
$\times$300 37.32 / 0.928 33.34 / 0.811 36.36 / 0.911 32.99 / 0.780 33.27 / 0.803 37.68 / 0.928 37.87 / 0.934 37.48 / 0.919 37.94 / 0.930
@@
$\times$200 41.70 / 0.912 41.16 / 0.885 43.43 / 0.954 40.86 / 0.884 40.26 / 0.885 42.25 / 0.924 44.85 / 0.969 43.91 / 0.968 45.11 / 0.971
@@ 4.1 Evaluation Protocols
Since PSNR and SSIM metrics are sensitive to overall image illumination, even minor shifts can lead to significant variations in these values. We apply the illumination correction method from the ELD paper [9]  [^1] See this part of the code. to the denoised images. When comparing our method with other state-of-the-art methods, we apply this illumination correction to all methods before evaluating PSNR and SSIM.
Note that the original ELD paper crops the image centers to a size of $512\times 512$ for PSNR/SSIM computation, whereas we compute them on the entire image at a resolution of $4240\times 2832$. In the following results, we apply the same measurement setup for all methods.
Noise Statistics. We also analyze the pixel-wise marginal distributions between the generated and real noise images in the test set, similar to the approach presented in the Noise Flow paper [13]. This evaluation prioritizes the model’s ability to capture the fundamental characteristics of the noise distribution. Specifically, we compute the histograms of real-world and generated noise and report the discrete Kullback-Leibler (KL) divergence between these histograms.
@@ 4.2 Implementation Details
We train our diffusion model using the SID Sony RAW image training set. For evaluation, in line with the approach used in the ELD paper [9], we evaluate our denoising network on a combination of the SID test and validation sets (referred to as the SID test set hereafter), using the same evaluation images as in the ELD paper. We also evaluate our model on the ELD Sony dataset.
To assess the quality of the synthetic data, we train denoising networks using the $L1$ loss between the network’s output and the reference clean image, then evaluate their denoising performance. For a fair comparison, we use the same U-Net architecture employed in [4].
Further implementation details are provided in the appendix.
Figure 4: Visual comparison of denoising results from networks trained on data synthesized using different methods, alongside comparisons with a model trained on real clean-noisy pairs from the SID training set. The first sample is from the SID test set, while the second is from the ELD test set. In the first example, our method produces cleaner results, while other methods leave some residual noise. The second example demonstrates that our method recovers text details more clearly and with fewer artifacts. Best viewed when zoomed in.
TABLE II: KL divergence comparisons of generated noise on the SID dataset of different data synthesis methods.
Poisson-Gaussian NoiseFlow PMN NoiseDiff
$\times$100 0.094 0.415 0.139 0.025
$\times$250 0.070 0.193 0.196 0.031
$\times$300 0.187 0.155 0.407 0.057
4.3 Noise statistic Comparison
To evaluate the quality of noise generated by different methods, we compare the KL divergence between the real noise distribution and the generated noise distributions of different data synthesis methods. The results are presented in Table II, excluding methods without released noise generation code. The results show that our method achieves the lowest KLD.